OpenStudy (anonymous):

How would you find the derivative of y = sin−1(6x + 1)?

myininaya (myininaya):

if \[y=\sin^{-1}(6x+1) => \sin(y)=6x+1 => y' \cos(y)=6 => y'=\frac{6}{\cos(y)}\] if sin(y)=(6x+1)/(1) (=opp/hyp) 

myininaya (myininaya):

\[\cos(y)=\frac{\sqrt{1^2-(6x+1)^2}}{1} (=\frac{adj}{hyp})\]

myininaya (myininaya):

\[y'=\frac{6}{\sqrt{1-(6x+1)^2}}\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

i tried applying that to another question and this is how far I got: \[y=25\arctan \sqrt{x}\] \[y=25\tan^{-1} (\sqrt{x})\] \[25\tan(y)=\left(\begin{matrix}\sqrt{x} \\ 25\end{matrix}\right)\] \[(\sec(y))(\tan(y))=\left(\begin{matrix}\sqrt{x} \\ 25\end{matrix}\right)\]

OpenStudy (anonymous):

what would i do next?

myininaya (myininaya):

\[y=25\tan^{-1}(\sqrt{x}) => \frac{y}{25}=\tan^{-1}(\sqrt{x})\] \[=>\tan(\frac{y}{25})=\sqrt{x}=\frac{\sqrt{x}}{1} (=\frac{opp}{adj})\]

myininaya (myininaya):